OpenStudy (baldymcgee6):

Find the secant line through 1^x at x=1/6 and x=1/3.

OpenStudy (baldymcgee6):

any ideas?

OpenStudy (baldymcgee6):

@hartnn

hartnn (hartnn):

u have any formulas for this?

OpenStudy (baldymcgee6):

no

hartnn (hartnn):

then let me search through net....

hartnn (hartnn):

1^x isn't 1 ?? can u post the pic of question or verify it just once?

OpenStudy (amistre64):

a secant line is one that passes thru a curve at 2 points

OpenStudy (amistre64):

instead of finding the equation of the tangent line at point "c"; you are just finding the equation of the line that passes thru the 2 points

hartnn (hartnn):

which 2 points 2 use? (1/6,0) and (1/3,0) ??

OpenStudy (amistre64):

the ones given yes;

OpenStudy (amistre64):

1/6, 1^(1/6) and 1/3,1^(1/3)

OpenStudy (baldymcgee6):

@amistre, isnt the function f(x) = 1^x just a line?

OpenStudy (baldymcgee6):

@amistre64

OpenStudy (baldymcgee6):

how can a secant line intercept a straight line at 2 different points?

hartnn (hartnn):

that was the sole reason i asked u to verify the question....

OpenStudy (baldymcgee6):

sorry i was gone, but that is the exact question

hartnn (hartnn):

u are right, y=1^x is same as line y=1 even with amistre's method, u will finally get equation of secant line as y=1 .....

OpenStudy (baldymcgee6):

yeah, doesnt make sense

OpenStudy (amistre64):

you have to ignore the geometric intepretation and just go with it as a line between 2 points. if we try to go by the geometric interpretation of a tangent for say: y=2x+1, we get y'= 2

OpenStudy (amistre64):

how do you determine the tangent to a line at a given point?

OpenStudy (baldymcgee6):

the derivative

OpenStudy (amistre64):

but by using the same "illogic" of a geometric definition. You cant have a tangent to a line anymore than you can have a secant to a line.

OpenStudy (baldymcgee6):

okay, so the question is a trick?

OpenStudy (amistre64):

its a play on words if anything. You have to take it into the abstract; a secant is a line between 2 specified points

OpenStudy (baldymcgee6):

okay, thanks for the help/clarification

OpenStudy (amistre64):

yw